Tom is twice as old as Dick. Dick is four years younger than Harry. Harry is eight years younger than Tom. How many years old is Harry?
Answer by mananth(16949)

Tom is twice as old as Dick.
Dick is four years younger than Harry.
Harry is eight years younger than Tom.
let tom be x years
Harry x-8
Dick (x-8)-4
Tom = 2 * Dick
x= 2((x-8)-4)
x=2x-16-8
-x=-24
x=24 Tom's age
Harry = 24-8=16 years